OpenStudy (anonymous):

log base 6 (x-1) + log base 6 (x+4)=2

OpenStudy (anonymous):

this means 6^2 = (x-1)(x+4) 36 = x^2+3x-4 x^2+3x-40 =0 x= 5 or -8

OpenStudy (anonymous):

careful!

OpenStudy (anonymous):

there is a small error here.

OpenStudy (anonymous):

ya thats what i thought..

OpenStudy (anonymous):

Nick Black Champion 0 log6(x−1)+log6(x+4)=2 log6(x−1)(x+4)=2 x2+3x−40=0 (x+8)(x−5)=0 x=-8,5

OpenStudy (anonymous):

you have to use anti logs

OpenStudy (anonymous):

where di u get 40?

OpenStudy (anonymous):

\[\log_6((x-1)(x+4))=2\] \[6^2=(x-1)(x+4)\] \[x^2+3x-4=36\] \[x^2+3x-40=0\] \[(x+8)(x-5)=0\]

OpenStudy (anonymous):

where is the error

OpenStudy (anonymous):

all this is good, but \[x=-8,x=5\] is wrong because you cannot take the log of a negative number so answer is only \[x=5\]

OpenStudy (anonymous):

thanks for the info

OpenStudy (anonymous):

no prolbem

OpenStudy (anonymous):

Yeah anti logarithms are greater than 0, so you have \[x - 1 > 0 , x+4>0\] Which results to \[x >1\] Therefore the solution x = -8 does not satisfy this condition. So x = 5 As satellite73 put.